OpenStudy (anonymous):

4 3/5 + 1 24/25 A. 6 14/25 B. 6 17/25 C. 5 14/25 D. 5 17/25

OpenStudy (solomonzelman):

Do you want to see it clearly? \[\huge\color{blue}{ 4\frac{3}{5} +1\frac{24}{25} } \color{red} { }\] \[\huge\color{blue}{ 4\frac{3\color{red} { \times 5 }}{5 \color{red} { \times 5 }} +1\frac{24}{25} } \] \[\huge\color{blue}{ 4\frac{15}{25} +1\frac{24}{25}}\] see what I did, can you add the whole numbers and fractions and put them together?

OpenStudy (anonymous):

Thanks :D

OpenStudy (solomonzelman):

Just tell me what your final answer is, so that we can make sure it's right, OK ?

OpenStudy (anonymous):

okay I got 6 14/25

OpenStudy (solomonzelman):

YES !!!! CORRECT !!!

OpenStudy (anonymous):

Thanks for your help !!!!!!!!! :D